### 14 Dec 2023 v1

# Training-free Zero-shot Composed Image Retrieval with Local Concept Reranking

Training-free Zero-shot Composed Image Retrieval with Local Concept Reranking
Shitong Sun
Queen Mary University of London
shitong.sun@qmul.ac.uk
   Fanghua Ye
University College London
fanghua.ye.19@ucl.ac.uk
   Shaogang Gong
Queen Mary University of London
s.gong@qmul.ac.uk
Composed image retrieval attempts to retrieve an image of interest from gallery images through a composed query of a reference image and its corresponding modified text. It has recently attracted attention due to the collaboration of information-rich images and concise language to precisely express the requirements of target images. Most of the existing composed image retrieval methods follow a supervised learning paradigm to perform training on a costly triplet dataset composed of a reference image, modified text, and a corresponding target image. To alleviate the demand for difficult-to-obtain labeled triplet data, recent methods have introduced zero-shot composed image retrieval (ZS-CIR), which aims to retrieve the target image without the supervision of human-labeled triplets but instead relies on image-text pairs or self-generated triplets. However, these methods are less computationally efficient due to the requirement of training and also less understandable, assuming that the interaction between image and text is conducted with implicit query embedding. In this work, we present a new Training-Free zero-shot Composed Image Retrieval (TFCIR) method which translates the query into explicit human-understandable text. This helps improve computation efficiency while maintaining the generalization of foundation models. Further, we introduce a Local Concept Reranking (LCR) mechanism to focus on discriminative local information extracted from the modified instruction. Extensive experiments on three ZS-CIR benchmarks show that the proposed approach can achieve comparable performances with state-of-the-art methods and significantly outperforms other training-free methods on the open domain datasets, CIRR and CIRCO, as well as the fashion domain dataset, FashionIQ.
Figure 1: A comparison of query processing between existing zero-shot composed image retrieval methods and ours. Top: Existing methods are fine-tuned on text-image pairs [25, 1] or task-specific triplets [10, 17]. Existing methods employ a fusion embedding for queries without explicit semantic meaning. Bottom: We propose a new method without the requirement of training. Our method processes the query with explicit semantics at both the global and local levels.
1 Introduction
Composed image retrieval (CIR) aims to retrieve the target image from gallery images through the composed query, which is specified in the form of an image plus some text that describes desired modifications to the input image [27, 3]. The text-image composed query can overcome the limitation of singular modality image retrieval, where the image is dense in semantics and text is sparse [7]. Specifically, text-image retrieval encounters challenges due to imprecise descriptions and an unlimited number of correct targets. For example, a text query `dog' can correspond to thousands of dog images in various breeds, poses, and environments. Meanwhile, image-image retrieval suffers from an expression limitation, lacking the ability to generalize to different visual contents [8]. This becomes infeasible when a single image cannot be tailored to users’ intentions. Traditional composed image retrieval methods [18] are primarily trained on datasets of triplets composed of a reference image, a relative caption, and a target image. However, these methods entail the expensive cost of collecting triplet datasets. To overcome this challenge, recent research has made efforts to consider zero-shot composed image retrieval (ZS-CIR).
Current methods for ZS-CIR are mainly based on the nearest neighbour search in pre-trained visual-language embedding spaces [24, 12]. These embeddings are further fine-tuned using either task-independent image-text pairs [25, 1] or additional task-specific image-text-image triplets [10, 17].
However, this direct usage has two limitations. Firstly, the efficiency of existing methods is compromised by training large foundation models on downstream image-text pairs or task-specific triplets. Secondly, current methods represent the composed query as an implicit holistic embedding, assuming that the interaction between the reference image and modified text is encoded. Typical interactions involve addition, negation, comparison, spatial relations variation, etc [1]. However, the interactions encoding image-image relations are typically not learned in foundation models' pretraining processes [24, 12], which are mainly trained on image-text pairs. Specifically, in image-text pairs, the text is an image caption with every word in the text having a correspondence in the target image. In contrast, in image-text-image triplets, the text is an instruction, encoding variations of the reference visual content and can be irrelevant to the target image.
To address the first problem of high computation, we have taken the lead to build a training-free baseline for zero-shot composed image retrieval. The baseline aims to convert the composed query, consisting of a reference image and modified text, into a single text query. It retrieves the target image in the aligned vision-language feature space [13] without the need for additional training. Utilizing the recently developed captioner BLIP2 [13] to translate the image to text modality and a large language model for text understanding, our mechanism is not only computationally efficient, being training-free, but also capable of generating human-understandable query text. However, querying with a single holistic embedding remains implicit and lacks explicit discriminative information from text-image interaction. To tackle this challenge, we transform implicit interaction between query text and image into explicit local text using a reranking mechanism. This mechanism not only enhances computational efficiency but also improves understandability. The comparison of query processing between conventional methods and our proposed methods is shown in Figure 1.
Our work is motivated by the recent development of vision-language models [15], which excel in local feature discrimination. However, properly converting implicit text-image queries to explicit text-only queries at the local level remains a challenge. The vision-language model [15] is designed to generate text but lacks the ability to rank multiple gallery images. To address this limitation, we introduce a scoring module based on the probability of text prediction. Specifically, we extract local discriminative text as a text understanding problem and force the vision-language model to judge whether the local attributes exist in the target images by answering `Yes' or `No'. Together with global ranking score of the query text and target images, we incorporate the probability of text prediction as reranking scores. With this local concept reranking mechanism, our model can generate human-understandable explicit attributes in the training-free framework. To summarize, the contributions of this work are as follows:
  • We are the first to propose a training-free method for zero-shot composed image retrieval (TFCIR), eliminating the need for training on either image-text pairs or task-specific triplets.
  • We introduce a two-stage approach for TFCIR which incorporates coarse global-level retrieval baseline (GRB) as the first stage and fine-grained local concept reranking (LCR) as the second stage. This multiple-criteria scoring mechanism enables the model to rank based on both global and local information.
  • Our method achieves comparable performances with state-of-the-art methods and obtains significant improvements over other training-free methods on three different datasets: CIRR, FashionIQ, and CIRCO.
Figure 2: An overview of the Training-Free Composed Image Retrieval (TFCIR) model which consists of: Top: Global Retrieval Baseline (GRB) transforms the text-image composed query into a text-only query with a Large Language Model (LLM)-generated pseudo target caption. The global score for retrieval is determined by the similarity between the pseudo target caption's text embedding and target image embeddings. Bottom: Local Concept Reranking (LCR) extracts discriminative local concepts using a task instruction prompt for Large Language Model. Based on the global score from GRB, the top-$K$ images are reranked using both the global and local scores, calculated on their discriminative local concepts. The local score is translated into text prediction probability, detecting the existence of local concepts by a vision-language model, i.e. LLaVA. Both the global and local retrieval processes are conducted within a visual-language aligned feature space, i.e. BLIP2.
2 Related Work
Composed Image Retrieval.
Composed image retrieval aims to retrieve the target image from a gallery set where the input query is in the form of more than one modality. A standard query is composed of an image and various interactions, including relative attributes [21], natural language [3, 27] or spatial layout [19], to describe the desired modifications. Among these modalities, natural language has gained attention as it serves as a common means of conveying intricate specifications between humans and computers. Consequently, composed image retrieval becomes almost indistinguishable from image retrieval queried by both image and modified text. The key challenge in this task lies in learning the comparison between images. Conventional composed image retrieval addresses this challenge by training on task-specific triplets datasets. However, the demand for triplets datasets in CIR has led to expensive costs. To alleviate the need for triplets data, recent methods explore zero-shot composed image retrieval without training on existing benchmarks. Among them, CASE [10] and TransAgg [17] construct additional task-specific triplets datasets, which lack potential scalability. Pic2Word [25] and SEARLE [1] avoid using task-specific triplets but perform fine-tuning on text-image pairs by converting reference images into a single token. All existing zero-shot composed image retrieval methods require a training stage and compromise the generalization of the existing foundation models. We take the lead in proposing a training-free method for zero-shot composed image retrieval.
Descriptive LLMs in Vision.
The advancements in foundation models open up new opportunities for understandable visual recognition. Recent works [30, 31, 20] have mainly focused on image classification tasks, generating class-level attributes from class labels using the prior knowledge of the language model. The key idea is to query the large language model to obtain a set of attributes that describe each class and further apply vision-language models [24] to classify images via LLM-generated attributes. These fine-grained attribute-based classifications inherently provide interpretability. However, all of the current works are based on class level and cannot be generalized to the broader topic of instance level. In this work, we extend the descriptive LLMs to generate instance-level attributes in a composed image retrieval task. Another relevant work in composed image retrieval CASE [10] provides the interpretability through visualization of gradient-based heatmaps on both image and text. However, this visualization can only know where the model looks instead of what the model sees. In contrast, we aim to provide concept-level attributes instead of pixel-level saliency.
Prompt Engineering.
With the rapid development of large foundation models and increasing training difficulty, the paradigm has gradually transferred from traditional supervised learning to more efficient prompt-based learning [16]. This approach involves applying various templates to model input and leveraging zero-shot or few-shot learning to transfer model knowledge to various downstream scenarios. This prompt engineering topic first originates from the natural language processing domain [2, 23] and is later introduced to the multimodality domain [33, 34]. Extensive prompts [9] have been explored including task instruction prompting [23, 6] for downstream tasks, in-context learning [5, 2] by showing examplars, chain-of-thought prompting [22, 28] for progressively leading to easier logic. In this work, we explore the task instruction prompting for composed image retrieval. Specifically, leveraging the power of language models, we explore one prompt for merging reference image captions and another prompt for extracting local concepts with randomly selected examples. This prompt engineering is designed to stimulate the reasoning ability of large language models.
3 Methodology
In this work, we introduce a training-free multi-level ranking mechanism for zero-shot composed image retrieval, leveraging language models to bridge the semantic gap between source and target images. The aim of the proposed method TFCIR is to preserve the generalization ability of the foundation model by being training-free, as well as enabling retrieval with both global features about holistic query information and local features about discriminative partial query information. Figure 2 shows an overview of the proposed method which includes a coarse global sentence-level retrieval as the first stage (Section 3.1) and a local concept-level reranking as the second stage (Section 3.2).
Let us define some notations at first. Given a reference image $I_{r}$ and modified text $T_{m}$ as input query, the aim of TFCIR is to retrieve the target image $I_{t}$ from the gallery set $\{I_{t}^{n}\}_{n=1}^{N}$ without training.
3.1 Global Retrieval Baseline
To retrieve images of interest from composed image-text queries in a training-free paradigm, we formulate a Global Retrieval Baseline (GRB) which converts the less-explored composed image retrieval task into the well-developed text-image retrieval task. Namely, we convert the text-image composed query into a text-only query and use existing aligned visual-language feature space [13] to retrieve the target image of interest. As shown in the upper part of Figure 2, the inference process of the baseline model consists of three steps, including reference image captioning, pseudo target caption generation and text-image retrieval through the aligned feature space.
For each query composed of a reference image $I_{r}$ and its modified text $T_{m}$, we leverage off-the-shelf captioner BLIP2 [13] to generate a brief caption $T_{r}$ of the reference image $I_{r}$ defined as Eq. (1). Further, instead of dealing with a multimodal fusion problem suffering from misalignment between instructive text and its reference image, we work on a language task with a requirement of reasoning. Specifically, we combine the generated reference image caption $T_{r}$ and the modified text $T_{m}$ by the Large Language Model (LLM) GPT4 by default to generate a pseudo target caption $T_{g}$ as global-level retrieval information in Eq. (2). The prompt $P_{g}$ to generate the global-level pseudo target caption is shown in the upper part of Figure 2.
$T_{r}=Captioner(I_{r}),$ (1)
$T_{g}=LLM(P_{g},T_{r},T_{m}).$ (2)
With the pseudo target caption, the global retrieval ranks target images by similarity of the text embedding of pseudo target caption and image embeddings of target images in the pre-learned visual-textual aligned feature space [13]. The global score for target image $I^{n}_{t}$ is calculated as follows:
$S_{g}^{n}=<E_{t}(T_{g}),E_{i}(I_{t}^{n})>,$ (3)
where <>absent<>< > is to calculate the cosine similarity, $E_{t}$ and $E_{i}$ are the text and image encoders of BLIP2 [13] respectively. For each query in this inference process, Eqs. (1)-(2) are conducted once and Eq. (3) is conducted $N$ times for all the target images. Consequently, the multimodal fusing issue has been converted into a text comprehension issue, and the visual-textual aligned space can be leveraged directly without training.
3.2 Local Concept Reranking
Even though the formulated baseline with an explicit semantic query is able to transfer a text-image composed query to a text-only query in a training-free method. It still differs from conventional text-image retrieval, which aligns an image and its corresponding caption, a description of the image content. However, in composed image retrieval, the comparison between images is encoded in the form of modification text, an instruction for adjusting the image including addition, negation, comparison, etc. In other words, every word in conventional text-image retrieval has its counterpart in the corresponding image and contributes to boosting the retrieval. In contrast, in composed image retrieval, not all words have a direct correspondence with the target image, particularly for negation of the original image and certain descriptions of the reference image. For example, `Change the dog to a cat' or `Remove the lemon' illustrate these scenarios, where `the dog' and `the lemon' should not appear in the target image. In light of this, identifying the local concepts that will actually appear in the target image instead of all concepts from the query, can provide discriminative information and reduce false positives.
Therefore, we introduce a Local Concept Reranking (LCR) module based on local discriminative concepts that should exist in target images to rerank the results from the global retrieval baseline. As depicted in the lower part of Figure 2, the LCR module consists of three steps: modification parsing, local concept re-identification, and scoring integration. This reranking process, incorporating explicit local features, enhances semantic information from a global level in GRB to a more fine-grained local level.
Modification Parsing.
To identify the local concepts that should exist in the target images, we parse the modified text $T_{m}$ by a large language model. We refer to these should-exist local concepts as positive concepts. By designing a local concept parsing prompt $P_{l}$ with manually crafted exemplars, we obtain the positive concepts in textual form. Notably, it becomes popular to leverage LLMs to generate text descriptions [30, 31, 20] to enhance visual perception. However, current descriptions are typically generated at the class level while our approach generates descriptions at the instance level, specifically tailored to each query. The process can be formulated as follows:
$T_{l}=LLM(P_{l},T_{m}),$ (4)
where $T_{l}$ denotes the positive local concepts in textual form. This modification parsing module can help identify positive concepts from the query, providing more discriminative semantic information.
Local Concept Re-identification.
Having obtained the positive concepts by parsing modified text in the query, we further determine whether the positive concepts exist for each target image. To detect the local concept existence, we utilize a multimodal model LLaVA [15], which is an instruction-following visual question-answering model, and returns the result by text. However, LLaVA can only detect the existence of local features and return a hard result with the text `Yes' or `No'. This hard result cannot be utilized in the ranking scenario directly. To address this issue, we unlock the last layer of the LLaVA and take the predicted probability of text `Yes' and `No' as the local score, which can be formulated as follows:
$Logit_{YES}^{n},Logit_{NO}^{n}=LLaVA(P_{re},T_{l},I^{n}_{t}),$ (5)
$S_{l}^{n}=softmax(Logit_{YES}^{n},Logit_{NO}^{n}),$ (6)
where $P_{re}$ is the prompt for LLaVA as shown in Figure 2 and $S_{l}^{n}$ is the local score for the $n$-th target image. To enhance efficiency and reduce computation, we only re-rank the top-$K$ results of the global baseline. This explicit comparison can help re-identify the fine-grained discriminative information and thus boost retrieval performance.
Score Integration Module.
Given the ranking scores from the first stage of the global retrieval baseline and the second stage of local concept reranking, we can achieve the final ranking score by simply summing the global score and local score. Namely, the final score $S^{n}$ for the $n$-th target image can be shown as follows:
$S^{n}=S_{g}^{n}+S_{l}^{n}.$ (7)
The final ranking is achieved by the $S^{n}$ value from maximum to minimum.
We summarize our proposed method in Algorithm 1.
Algorithm 1 Training-Free Zero-Shot Composed Image Retrieval with Local Concept Reranking.
1:Input: Query of reference image $I_{r}$ and modified text $T_{m}$. Gallery set of target images {$I_{t}\}_{n=1}^{N}$. Hyperparameter rerank top $K$.
2:Global Retrieval Baseline:
3:Generate reference image caption $T_{r}$ via Eq. (1).
4:Generate pseudo target caption at the global level $T_{g}$ via Eq. (2).
5:Calculate global score $S^{n}_{g}$ via Eq. (3) and achieve top-$K$ ranking results from maximum to minimum.
6:Local Concept Reranking:
7:Modification parsing to extract local positive concepts $T_{l}$ via Eq. (4).
8:for each target image $n$ = 1 to $K$ in the top-$K$ ranking results returned by the global retrieval baseline do
9:     Calculate local score $S_{l}^{n}$ via Eq. (5), Eq. (6).
10:end for
11:Calculate final score $S^{n}$ for target images via Eq. (7).
4 Experiments
4.1 Datasets and Evaluation Metrics
Recall@$K$ Recall@${}_{\text{Subset}}@K$
Training data Method $K=1$ $K=5$ $K=10$ $K=50$ $K=1$ $K=2$ $K=3$
Triplets CASE  [10] 35.40 65.78 78.53 94.63 64.29 82.66 91.61
TransAgg-combined [17] 37.87 68.88 - 93.86 69.79 - -
TransAgg-template [17] 38.10 68.42 - 93.51 70.34 - -
Text-image pairs PALAVRA [4] 16.62 43.49 58.51 83.95 41.61 65.30 80.94
Pic2Word [25] 23.90 51.70 65.30 87.80 - - -
SEARLE [1] 24.87 52.31 66.29 88.58 53.80 74.31 86.94
Training-free Text-only-M 21.49 45.78 57.11 81.33 64.80 82.72 91.78
Text-only-C 13.66 37.78 51.01 80.60 42.65 66.77 82.87
LLaVA-CIR 11.90 29.21 39.74 62.96 37.71 57.52 74.43
GRB 24.19 52.07 65.48 85.28 60.17 79.13 89.34
GRB + LCR 30.92 56.99 68.58 85.28 66.67 78.68 82.60
Table 1: Quantitative results on CIRR test set. Best scores are highlighted in bold.
We conduct experiments on three publicly available CIR benchmarks, CIRR [18], FashionIQ [29] and CIRCO [1]. All datasets are in the form of triplets, composed of a reference image, modified text, and a target image. FashionIQ, focused on the fashion domain, contains garment images that can be divided into three categories: dress, shirt and toptee. Each image in FashionIQ contains a single subject positioned centrally with a clean background. CIRR is composed of real-life images extracted from NLVR2 [26], which contains rich visual content against various backgrounds. CIRCO, an open-domain dataset, includes 1020 queries divided into 220 for the validation set and 800 for the test set. The dataset utilizes all 120K images of COCO [14] as the gallery set and provides multiple ground truths for each query. Notably, we adopt zero-shot evaluation to test on the test sets of CIRR and CIRCO, and the validation set of FashionIQ following the standard evaluation protocol [25, 1]. The compared conventional methods are fine-tuned on task-specific datasets but our method is training-free. For single ground truth datasets CIRR and FashionIQ, we utilize the top K recall metric for evaluation, and for the multiple ground truth dataset CIRCO, we utilize mean Average Precision (mAP). Specifically, CIRR is evaluated on both the entire gallery and a subset gallery respectively. The subset gallery consists of five highly similar images for each query. FashionIQ evaluation is conducted separately for each category following previous works [25, 1].
4.2 Implementation Details
We leverage BLIP2 with OPT-2.7b [32] as the default captioner and its image-text matching module with aligned feature space to retrieve target images. For LLM supplying context knowledge, we leverage ChatGPT4-turbo by default to generate pseudo target captions or local discriminative attributes (concepts). For the generated text, we employ a cleaning process to standardize the format and enhance the quality of the text. To achieve scores corresponding to local concepts, we utilize LLaVA-v1.5-13b by default and rerank top 50 target images. All the experiments are conducted on NVIDIA A100 GPUs.
Shirt Dress Toptee Average
Training data Method R@$10$ R@$50$ R@$10$ R@$50$ R@$10$ R@$50$ R@$10$ R@$50$
Triplets TransAgg-Combined [17] - - - - - - 34.36 55.13
Text-image pairs PALAVRA [4] 21.49 37.05 17.25 35.94 20.55 38.76 19.76 37.25
Pic2Word [25] 26.20 43.60 20.00 40.20 27.90 47.40 24.70 43.70
SEARLE-XL-OTI [1] 30.37 47.49 21.57 44.47 30.90 51.76 27.61 47.90
Training-free Text-only-M 25.76 41.61 18.52 35.94 27.59 46.91 23.93 41.44
Text-only-C 21.64 39.60 11.80 27.42 20.14 38.25 17.85 35.07
LLaVA-CIR 9.32 19.33 6.15 14.72 9.48 19.22 8.32 17.75
GRB 34.54 55.15 24.14 45.56 33.55 53.60 30.74 51.44
Table 2: Quantitative results on FashionIQ validation set. Best scores are highlighted in bold.
mAP@$K$
Training data Method $K=5$ $K=10$ $K=25$ $K=50$
Text-image pairs PALAVRA [4] 4.61 5.32 6.33 6.80
Pic2Word [25] 8.72 9.51 10.64 11.29
SEARLE [1] 11.68 12.73 14.33 15.12
Training-free GRB 23.76 25.31 28.19 29.17
GRB + LCR 25.38 26.93 29.82 30.74
Table 3: Quantitative results on CIRCO test set. Best scores are highlighted in bold.
Figure 3: Qualitative results on CIRR test set. The GRB is the global retrieval baseline, which can merge holistic information from reference image caption and modified text. LCR is local concept reranking, which can extract discriminative local information to rerank based on the results of the global retrieval baseline.
CIRR
Method Captioner LLM R@$1$ R@$5$ R@$10$ R@$50$
GRB BLIP2 GPT4-turbo 24.19 52.07 65.48 85.28
GRB BLIP2 GPT3.5 21.49 45.78 57.11 81.33
GRB instructBLIP GPT4-turbo 22.60 49.88 62.55 84.99
GRB instructBLIP GPT3.5 20.75 45.54 58.10 79.42
Table 4: Ablation studies on CIRR test set. The best scores are highlighted in bold.
CIRR
Method Prompt $P_{l}$ R@$1$ R@$5$ R@$10$
GRB + LCR default 30.92 56.99 68.58
GRB + LCR w/o exemplar 29.54 58.17 69.86
GRB + LCR w/o negative 30.39 58.87 70.27
Table 5: Prompt $P_{l}$ variations with the result on CIRR test set. $P_{l}$ is aimed to parse the modification text and extract local positive concept.
4.3 Comparison with the SOTAs
Competitors.
We compare the proposed TFCIR with several state-of-the-art zero-shot composed image retrieval methods classified into three groups:
  • Trained on text-image pairs: Pic2Word [25], SEARLE [1] and PALAVRA [4]. All of them are textual inversion modules based on the foundation model CLIP [24]. They convert each image to a pseudo word token by fine-tuning on text-image pairs and retrieve in the CLIP pre-learned feature space. PALAVRA is aimed for the personalized image retrieval task with ZS-CIR results reported from [1]. SEARLE [1] additionally includes GPT-powered regularization and knowledge distillation.
  • Trained on task-specific triplets: CASE [10] and TransAgg [17]. They both propose new task-specific triplets, namely LaSCo for CASE and Laion-CIR for TransAgg. Both of them implement specific layers from the existing foundation model BLIP [12].
  • Training-free methods: We implement three training-free methods for comparison. Text-only-M conducts retrieval directly with the modified text of the triplets. Text-only-C performs retrieval directly with {reference caption} + `change to' + {modified text}, i.e., the concatenation of the modified text and the reference caption of the reference image, which is achieved from the BLIP2 captioner. LLaVA-CIR is achieved by inputting modified text and image, instructing the LLaVA model to output the target caption directly. All of these three methods calculate the cosine similarity distance in the aligned feature space of BLIP2 [13].
4.4 Main Results
CIRR.
The experimental results for the CIRR test set are presented in Table 1. As observed, our baseline method GRB, employing global-level retrieval achieves comparable performance with state-of-the-art ZS-CIR methods SEARLE [1] and Pic2Word [25]. Notably, both of them are trained on text-image pairs, while our method is training-free for saving computational resources. Furthermore, with reranking module LCR incorporating discriminative local concepts, our method gains a significant improvement of 6.72% in Recall@1. Local concepts provide discriminative information, leading to noticeable enhancements. In comparison with three other training-free methods, our method GRB + LCR has an average improvement of 9.01% over Text-only-M, 14.68% over Text-only-C, and 24.49% over LLaVA-CIR respectively. Among these training-free methods, LLaVA-CIR processes detailed information from the reference image, Text-only-C considers the full caption of the reference image, and Text-only-M lacks input from the reference image. It is evident that directly implementing reference image information without careful selection can distract the model and compromise performance. In contrast, the modified text itself provides strong discriminative information and yields competitive results. Our method GRB selectively extracts information from reference image captions by prompt-enhanced language model, resulting in improved performance. Notably, our proposed baseline model can outperform existing training-needed methods [1, 25] by an average improvement of 2.43% over SEARLE and 3.27% over Pic2word. For the evaluation on subset gallery, the discriminative ability is required. Our LCR provides a strong score for the target local concept, increasing the score difference between images, and resulting in a significant improvement of 6.5% over the baseline GRB.
FashionIQ.
We present the results on the FashionIQ validation set in Table 2. We can observe that our GRB baseline alone achieves better performance than the existing state-of-the-art ZS-CIR SEARLE, with an average improvement of 6.81% in Recall@5 and 3.54% in Recall@10. FashionIQ is a fine-grained dataset with similar clean backgrounds and fine-grained variations. We can observe that modified text alone provides strong discriminative information. With the increase of reference image information, the model gets distracted and leads to lower performance from Text-only-C to LLaVA-CIR.
CIRCO.
As shown in Table 3, we present the performance on the CIRCO test set. As an open domain dataset with multiple ground truths from a much larger gallery set, CIRCO imposes higher requirements of fine-grained discriminative ability. Both of our methods show obvious improvements over existing methods. Specifically, our baseline with global-level retrieval GRB outperforms SEARLE by 12.08% in mAP@5. Additionally, reranking with local concepts LCR further yields over 1% average improvement. The substantial enhancement observed in CIRCO underscores the superiority of addressing reasoning in text over aligning the two modalities.
Visualization Results.
The visualization of the proposed baseline method GRB and reranking method LCR is shown in Figure 3. The baseline method GRB provides coarse direction and the reranking method LCR can identify local discriminative concepts.
4.5 Ablation Studies
Evaluating Variants of Captioner and LLMs.
In this section, variations of captioner and large language models are identified to provide context knowledge. We implement instructBLIP [11] to generate a descriptive caption by using the prompt `Describe the content of the image in detail'. Additionally, different versions of ChatGPT are deployed to explore the impact of the large language models on the reasoning process. As shown in Table 4, we observe that instructBLIP provides detailed reference information but misleads the model, which is aligned with the observation in Sec. 4.4. Meanwhile, the large language model GPT-4 outperforms GPT3.5, indicative of enhanced reasoning capabilities.
Evaluating Variants of Prompts.
To further evaluate the sensitivity of our model to prompts, we conducted experiments employing diverse prompts, incorporating variations in the number of exemplars and the omission of the negative category. The outcomes of these experiments are summarized in Table 5. Specifically, adhering to the default prompt denoted as $P_{l}$ in Figure 2, we conducted experiments involving the removal of exemplars and the exclusion of the negative category respectively. Under the default prompt $P_{l}$ with manual exemplars and two categories of local concepts, namely positive and negative, the model exhibits superior performance in terms of Recall@1 and a decrease in Recall@5 and Recall@10. In the broader context, our model demonstrates robust performance against prompt variations.
Limitations and Future Work.
From the experimental results, it becomes evident that a judicious selection process is imperative when incorporating information from reference images. In this study, we address this by employing a two-step approach involving initial captioning, followed by the utilization of a Large Language Model (LLM) to generate the pseudo-target caption. The selection of reference information remains an avenue deserving further investigation.
5 Conclusion
This paper presents an innovative training-free method for zero-shot composed image retrieval, incorporating reranking strategy informed by local concepts. Initially, a global retrieval baseline is established by transforming the text-image composed query into a text-only query. Subsequently, we enhance the retrieval process by extracting local concepts through modified parse techniques and implementing reranking based on the outcomes obtained in the initial stage. The proposed method demonstrates both high efficiency and competitive performance when compared with existing zero-shot composed image retrieval (ZS-CIR) methods.